# Two-level Nonlinear Preconditioning Methods for Flood Models Posed on Perforated Domains

### v1

@@ Abstract
This article focuses on the numerical solution of the Diffusive Wave equation posed in a domain containing a large number of polygonal perforations. These numerous perforations represent structures in urban areas and this model problem is used to model urban floods. This article relies on the work done in a previous article by the same authors, in which we introduced low-dimensional coarse approximation space for the linear Poisson equation based on a coarse polygonal partitioning of the domain. Similarly to other multi-scale numerical methods, this coarse space is spanned by locally discrete harmonic basis functions. We show that this coarse space also lends itself to the linearized diffusive wave problem that is obtained at each iteration of Newton’s method. Furthermore, we present nonlinear preconditioning techniques, including a two-level RASPEN and Two-step method, which can significantly reduce the number of iterations compared to the traditional Newton’s method. Numerical examples illustrating the performance of the proposed methods include a large-scale test case based on topographical data from the city of Nice.
journal: Journal TBD
@@ 1 Introduction
In this article, we apply domain decomposition and multi-scale approaches to solve nonlinear equations, both stationary and time-dependent. The main focus of the paper is the Diffusive Wave model [2], given by
@@
where $z_{b}(\mathbf{x})$ represents bathymetry/topography, $u$ is total water level, $h(u,z_{b}(\mathbf{x}))=\max(u-z_{b}(\mathbf{x},0)$ is the depth of the water with respect to bathymetry, $c_{f}$ denotes friction, $\gamma\leq 1$, and $1\leq\alpha\leq 2$. Generally, we take the source term $f$ as constant; this source term can be considered to represent rainfall. The Diffusive Wave Equation (DWE) is doubly nonlinear in the sense that there are nonlinearities both in the $(u-z)$ term and in the gradient term. Additionally, the equation becomes degenerate, as $\kappa(u,\nabla u)\rightarrow 0$ as $h\rightarrow 0$ and $\kappa(u,\nabla u)\rightarrow\infty$ as $\nabla u\rightarrow 0$. We can think of the DWE as a simplification of the 2D Shallow Water Equations with zero-inertia assumptions ; for this reason, the DWE is often referred to as the Zero Inertia (ZI) equations [2]. This equation is commonly used to model overland flows.
@@
With this, let $D$ be an open simply connected polygonal domain in $\mathbb{R}^{2}$. We denote by $\left(\Omega_{S,k}\right)_{k}$ a finite family of perforations in $D$ such that each $\Omega_{S,k}$ is an open connected polygonal subdomain of $D$. The perforations are mutually disjoint, that is $\overline{\Omega_{S,k}}\cap\overline{\Omega_{S,l}}=\emptyset$ for any $k\neq l$. We denote $\Omega_{S}=\bigcup_{k}\Omega_{S,k}$ and $\Omega=D\setminus\overline{\Omega_{S}}$, assuming that the family $\left(\Omega_{S,k}\right)_{k}$ is such that $\Omega$ is connected. The model problem of focus is given by the following: for a fixed time interval $(0,T]$,
$\left\{\begin{array}[]{rll}\displaystyle\partial_{t}u-\text{div}\left(\kappa(u,\nabla u)\nabla u\right)&=&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0}\qquad\mbox{in}\qquad\Omega\times(0,T],\\ \displaystyle u&=&g\qquad\mbox{on}\qquad(\partial\Omega\setminus\partial\Omega_{S})\times(0,T],\\ \displaystyle{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\kappa(u,\nabla u)}\frac{\partial u}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad(\partial\Omega\cap\partial\Omega_{S})\times(0,T],\\ \end{array}\right.$ (1)
@@
Depending on the geometrical complexity of the computational domain, the numerical solution of the PDE may become increasingly challenging. Our goal is to numerically solve the model problem on these perforated domains by using domain decomposition techniques. The latter use a divide and conquer approach to partition the domain into multiple smaller subdomains. Therefore, one has two “levels” of space discretization; the first level is based on a coarse polygonal partitioning of $\Omega$, while the second one is associated with the fine-scale triangulation and is designed to resolve the small-scale details of the model domain. The fine-scale triangulation is generated by Triangle [27]. Domain decomposition (DD) approaches to solve nonlinear PDEs can generally split into two categories: Newton-Krylov Schwarz (NKS) and Schwarz Newton-Krylov (SNK) methods.
@@
In NKS methods [6], the problem is first linearized, typically by Newton’s method or an inexact Newton method, which solve a linear system at each iteration. It is well-known that for large/complex linear systems, a direct linear solve can become quite expensive and time-consuming. In this case, one can then apply typical linear DD methods to this linear Newton system; i.e., use standard domain decomposition preconditioners coupled to a Krylov on the linear system to be solved at each Newton iteration. Therefore, in the case of NKS methods applied to a nonlinear problem $F(\mathbf{u})=0$, the system remains unchanged, and we use DD methods to make the linear solve more efficient.
@@
Stiff nonlinearities arise in many nonlinear problems, including problems heterogeneous problems with high-contrast coefficients. These stiff nonlinearities can cause a plateau in the norm of the Newton residual before/if the quadratic region of convergence is reached. In these cases in a global Newton’s method, one may be forced to take very small time steps, or damp the Newton iteration to an extremely small damping factor. To combat this, more recently, SNK methods were introduced; these methods do not begin with a linearization and are commonly referred to as nonlinear preconditioning methods. In SNK methods, the nonlinear system itself is changed, unlike in NKS methods. The nonlinear preconditioners involve the solution of local nonlinear problems and can therefore localize stiff nonlinearities.
@@
Specifically, given a nonlinear system $F(\mathbf{u})=0$, these SNK preconditioning methods solve a new nonlinear system $\mathcal{F}(\mathbf{u})$ via Newton’s or inexact Newton’s method, where if $\mathcal{F}(\mathbf{u})=0$, then $F(\mathbf{u})=0$ (i.e. the new nonlinear system shares solutions with the original nonlinear system). This system $\mathcal{F}(\mathbf{u})=0$ is referred to as the nonlinearly preconditioned system. The latter is generated by applying Newton’s or inexact Newton’s method to a fixed point equation. The goal is for the nonlinearly preconditioned system to be solved faster than the original one; that is, to more quickly advance to the region of quadratic convergence for Newton’s method and reduce the number of exact or inexact Newton iterations.
@@
In the Additive Schwarz Preconditioned Inexact Newton (ASPIN) [7], the authors introduce a method, where the nonlinearly preconditioned system $\mathcal{F}(u)=0$ consists of more balanced/uniform nonlinearities. ASPIN involves the solution of local nonlinear problems, using similar restriction and extension operators that are used in the linear case. Then, a global linear problem is solved via inexact Newton. In theory, for difficult problems with intense nonlinearities, this technique can accelerate convergence when compared to a typical Newton or inexact Newton method. Additionally, it can aid in the occurrence of stagnation in Newton’s method and desensitize Newton to the initial guess. We remark as well the existence of the MSPIN method [24], where generally, the partitioning of degrees of freedom is based on field type (field-splitting) instead of the typical subdomain partitioning. A main difference between the methods is that MSPIN can be thought of as analogous to the Multiplicative Schwarz (MS) method for linear systems, a generally non-parallel method, while ASPIN is analogous to the Additive Schwarz method (ASM). With this, one can also think of ASPIN and MSPIN as varying from each other in a similar manner to Jacobi and Gauss-Seidel methods. In MSPIN and ASPIN, inexact Newton solves are computed for the nonlinearly preconditioned system.
@@
The authors of [10] introduced SRASPEN, which is obtained in a similar way by applying Newton’s method to the fixed point equation resulting from the nonlinear SRAS (Substructured Restricted Additive Schwarz) method. Unlike the aforementioned methods, the SRASPEN method works on nonoverlapping subdomains; these nonoverlapping domain decomposition methods are often referred to as substructuring methods. Within the same spirit, the authors of [11] introduced DNPEN, a nonlinear preconditioning method which applies Newton’s method to the Dirichlet-Neumann (DN) method, which is a substructuring method based on matching fluxes along subdomain interfaces. Additionally, the authors of [25] propose an adaptive nonlinear preconditioning technique which “turns off” the nonlinear preconditioning for outer Newtons where it is not necessary.
@@
The aforementioned techniques can be considered as being left preconditioning methods, as they are solving a new tangential system $\mathcal{F}(\mathbf{u})=0$ by exact or inexact Newton. We also remark that there are a family of right nonlinear preconditioning techniques; these methods change the unknowns of the original system. We can give as an example the following works [20, 26, 21, 22, 23], where the authors introduced nonlinear preconditioners analogous to the Finite Element Tearing and Interconnecting (FETI), Finite Element Tearing and Interconnecting Dual-Primal (FETI-DP), and Balancing Domain Decomposition by Constraints (BDDC) methods. Additionally, the authors of [8] introduced the Nonlinear Elimination Preconditioned Inexact Newton (NEPIN), where “problem” components of the residual are implicitly eliminated and a resulting modified Newton direction is computed. However, the focus of this article is left nonlinear preconditioning techniques based on overlapping subdomains.
@@
As is the case for linear DD methods, two-level methods (methods which include a coarse correction/level), are often needed for scalability in nonlinear preconditioning methods. Particularly, the number of GMRES iterations for the global tangential solve will not be scalable without a coarse correction. The coarse component allows for global communication across all subdomains. In the original article introducing RASPEN [13], the authors proposed a Full Approximation Scheme (FAS) to apply the coarse level. For MSPIN, a multiplicative coarse correction counterpart was introduced. In [17], the author propose a linear coarse level for ASPIN and in [15], the authors provide numerical evidence of scalability of two-level variants of ASPIN and RASPEN. In particular, they propose multiple approaches of adding the coarse space correction (both additively and multiplicatively) based on Galerkin projections, and a numerical comparison of the methods is reported for different types of coarse spaces.
@@
In this paper, we rely on the work done in [4], in which we introduced an MsFEM coarse space for the linear Poisson problem posed on the perforated domains. Here, for nonlinear problems, we focus on left nonlinear preconditioning based on overlapping subdomains, coupled to techniques that use the coarse space proposed in [4] to provide scalability. We propose this linear coarse space as a component of a two-level Restricted Additive Schwarz preconditioner for a linearized Newton system which arises from the linearization of a nonlinear equation. Additionally, we use this coarse space as a component for multiple nonlinear methods including a two-level RASPEN method and a “Two-step” method originally introduced in [9] under the name NKS-RAS. While the coarse space of [4] was originally created based on the linear Poisson equation, we find that it is a fitting coarse space for nonlinear problems as well, and we use this in the later-mentioned two-level nonlinear preconditioning methods. For these nonlinear methods applied to a time-dependent problem, we implement an adaptive local time-stepping method which allows us to avoid a global time step reduction. In addition, this article serves as a numerical application of nonlinear preconditioning methods applied to nonlinear problems, containing multiple numerical examples with the main example being the doubly nonlinear Diffusive Wave model posed on complex domains containing multiple perforations. For all proposed methods, we provide the detailed complexities/costs of each iteration.
@@
The remainder of the paper is structured as follows. Section 2 provides a detailed description of the finite-element/ finite-volume hybrid discretization used for the Diffusive Wave Equation. Section 3 describes the coarse space from [4], which will be used in the two-level methods described in the article. Section 4 summarizes domain decomposition methods for linear problems, presenting the well-known linear Additive Schwarz and Restricted Additive Schwarz methods. Section 5 describes the proposed methods in detail, followed by Section 5.1 which discusses the cost/complexity of the proposed algorithms. Section 6 provides numerical results, with experiments based on a porous media equation posed on both an L-shaped domain and a large urban domain, followed by the Diffusive Wave equation posed on a large urban domain. Section 7 concludes with a summary and description of future work.
2 2D spatial discretization of Diffusive Wave equation
@@
We now proceed with a discretization of the Diffusive Wave model (1). Let $T_{f}>0$ be the final flow simulation time and let $0=t_{0}<t_{1}<\ldots<t_{N}=T_{f}$ be a family of real numbers such that $\Delta t_{n}=t_{n+1}-t_{n}$.
@@
$\left\{\begin{array}[]{rll}\displaystyle\frac{u^{n+1}-u^{n}}{\Delta t_{n}}-\text{div}\left(c_{f}\kappa(u^{n+1},\nabla u^{n})\nabla u^{n+1}\right)&=&0\qquad\mbox{in}\qquad\Omega,\\ \displaystyle u^{n+1}&=&g\qquad\mbox{on}\qquad\partial\Omega\setminus\partial\Omega_{S},\\ \displaystyle{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c_{f}\kappa(u^{n+1},\nabla u^{n})}\frac{\partial u^{n+1}}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad\partial\Omega\cap\partial\Omega_{S},\end{array}\right.$ (2)
@@
Let $(\mathbf{x}_{i})_{i=1}^{N_{\Omega}}$ denote the set of triangulation points; the set of point indices $\{1,2,...,N_{\Omega}\}$ is denoted by $\mathcal{N}$. We denote the nodal “hat” basis functions $(\eta^{\ell})_{\ell=1}^{N_{\Omega}}$. Furthermore, we define $u_{i}^{n+1}$ and $u_{\ell}^{n+1}$ as the nodal values of the discrete solution at nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{\ell}$ at $t_{n+1}$, respectively. Additionally, we set ${\cal N}_{T}=\{i\in{\cal N}\,\,|\,\,\mathbf{x}_{i}\in\overline{T}\}$ for any triangle $T\in{\cal T}$ ($T$ assumed to be open) and denote $\mathcal{T}_{i}$ as the subset of triangles connected to node $i$ such that $\mathcal{T}_{i}=\{T\in\mathcal{T}\,\,|\,\,\mathbf{x}_{i}\in\overline{T}\}$. Furthermore, we denote by ${\cal N}_{i}$ the set of notes adjacent to node $i$ such that ${\cal N}_{i}=\bigcup_{T\in\mathcal{T}_{i}}{\cal N}_{T}.$ Finally, let ${\cal N}_{D}$ denote the set of Dirichlet nodes such that ${\cal N}_{D}=\{i\in{\cal N}\,\,|\,\,\mathbf{x}_{i}\in\partial\Omega\setminus\Omega_{S}\}$.
In order to enhance the stability of the numerical scheme, we modify the finite element scheme by means of discretization techniques from finite volume methods. In particular, we introduce mass lumping in the accumulation and upwinding in diffusion term. This discretization technique is generally referred to as a Finite Volume- Finite Element (FV-FE) discretization [12], or it is often referred to as a Control Volume Finite Element method.
@@
$\tau_{i\ell,T}^{n}=-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c_{f}}\Bigl{|}\nabla u_{h}^{n}|_{T}\Bigr{|}^{1-\gamma}\int_{T}\nabla\eta_{\ell}\cdot\nabla\eta_{i}\,{\rm d}\mathbf{x}\qquad\mbox{and}\qquad\tau_{i\ell}^{n}=\sum_{T\in{\cal T}_{i}\cap{\cal T}_{\ell}}\tau_{i\ell,T}^{n},$

### v2

@@ Abstract
This article focuses on the numerical solution of the Diffusive Wave equation posed in a domain containing a large number of polygonal perforations. These numerous perforations represent structures in urban areas, and this problem is used to model urban floods. This article relies on the work done in a previous article by the same authors, in which we introduced low-dimensional coarse approximation space for the linear Poisson equation based on a coarse polygonal partitioning of the domain. Similarly to other multi-scale numerical methods, this coarse space is spanned by locally discrete harmonic basis functions. We show that this coarse space also lends itself to the linearized Diffusive Wave problem that is obtained at each iteration of Newton’s method. Furthermore, we present nonlinear preconditioning techniques, including a Two-level RASPEN and Two-step method, which can significantly reduce the number of iterations compared to the traditional Newton’s method. For all proposed methods, we provide a discussion regarding the complexity of the algorithms. Numerical examples illustrating the performance of the proposed methods include a large-scale test case based on topographical data from the city of Nice.
journal: Computers and Mathematics with Applications
@@ 1 Introduction
In this article, we apply domain decomposition and multi-scale approaches to solve stationary and time-dependent partial differential equations. The main focus of the paper is the Diffusive Wave model [29, 2], given by
@@
where $z_{b}(\mathbf{x})$ represents bathymetry/topography, $u$ is total water level, $h(u,z_{b}(\mathbf{x}))=\max(u-z_{b}(\mathbf{x},0)$ is the depth of the water with respect to bathymetry, $c_{f}$ denotes friction, $\gamma\leq 1$, and $1\leq\alpha\leq 2$. Generally, we take the source term $f$ as constant; this source term can be considered to represent rainfall. The Diffusive Wave Equation (DWE) is doubly nonlinear in the sense that there are nonlinearities both in the $(u-z_{b})$ term and in the gradient term. Additionally, the equation becomes degenerate, as $\kappa(u,\nabla u)\rightarrow 0$ as $h\rightarrow 0$ and $\kappa(u,\nabla u)\rightarrow\infty$ as $\nabla u\rightarrow 0$. We can think of the DWE as a simplification of the 2D Shallow Water Equations with zero-inertia assumptions ; for this reason, the DWE is often referred to as the Zero Inertia (ZI) equations [2]. This equation is commonly used to model overland flows.
@@
With this, let $D$ be an open, simply connected polygonal domain in $\mathbb{R}^{2}$. We denote by $\left(\Omega_{S,k}\right)_{k}$ a finite family of perforations in $D$ such that each $\Omega_{S,k}$ is an open connected polygonal subdomain of $D$. The perforations are mutually disjoint, that is $\overline{\Omega_{S,k}}\cap\overline{\Omega_{S,l}}=\emptyset$ for any $k\neq l$. We denote $\Omega_{S}=\bigcup_{k}\Omega_{S,k}$ and $\Omega=D\setminus\overline{\Omega_{S}}$, assuming that the family $\left(\Omega_{S,k}\right)_{k}$ is such that $\Omega$ is connected. The model problem of focus is given by the following: for a fixed time interval $(0,T]$,
$\left\{\begin{array}[]{rll}\displaystyle\partial_{t}u-\text{div}\left(\kappa(u,\nabla u)\nabla u\right)&=&0\qquad\mbox{in}\qquad\Omega\times(0,T],\\ \displaystyle u&=&g\qquad\mbox{on}\qquad(\partial\Omega\setminus\partial\Omega_{S})\times(0,T],\\ \displaystyle\kappa(u,\nabla u)\frac{\partial u}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad(\partial\Omega\cap\partial\Omega_{S})\times(0,T],\\ \end{array}\right.$ (1)
@@
Depending on the geometrical complexity of the computational domain, the numerical solution of the PDE may become increasingly challenging. Our goal is to numerically solve the model problem on these perforated domains by using domain decomposition techniques. The latter use a divide and conquer approach to partition the domain into multiple smaller subdomains. Therefore, one has two “levels” of space discretization; the first level is based on a coarse polygonal partitioning of $\Omega$, while the second one is associated with the fine-scale triangulation and is designed to resolve the small-scale details of the model domain. The fine-scale triangulation is generated by Triangle [30].
@@
Domain decomposition (DD) approaches to solve nonlinear PDEs can generally split into two categories: Newton-Krylov Schwarz (NKS) and Schwarz Newton-Krylov (SNK) methods. In NKS methods [6], the problem is first linearized, typically by Newton’s method or an inexact Newton method, which solve a linear system at each iteration. It is well-known that for large/complex linear systems, a direct linear solve can become quite expensive and time-consuming. In this case, one can then apply typical linear DD methods to this linear Newton system; i.e., use standard domain decomposition preconditioners coupled to a Krylov method on the linear system to be solved at each Newton iteration. Therefore, in the case of NKS methods applied to a nonlinear problem $F(\mathbf{u})=0$, the system remains unchanged, and we use DD methods to make the linear solve more efficient.
@@
Stiff nonlinearities arise in many nonlinear problems, including problems such as heterogeneous problems with high-contrast coefficients. These stiff nonlinearities can cause a plateau in the norm of the Newton residual before/if the quadratic region of convergence is reached. In these cases in a global Newton’s method, one may be forced to take very small time steps, or damp the Newton iteration to an extremely small damping factor. To combat this, more recently, SNK methods were introduced; these methods do not begin with a linearization and are commonly referred to as nonlinear preconditioning methods. In SNK methods, the nonlinear system itself is changed, unlike in NKS methods. The nonlinear preconditioners involve the solution of local nonlinear problems and can therefore localize stiff nonlinearities.
@@
Specifically, given a nonlinear system $F(\mathbf{u})=0$, these SNK preconditioning methods solve a new nonlinear system $\mathcal{F}(\mathbf{u})$ via Newton’s or inexact Newton’s method, where if $\mathcal{F}(\mathbf{u})=0$, then $F(\mathbf{u})=0$ (i.e., the new nonlinear system shares solutions with the original nonlinear system). This system $\mathcal{F}(\mathbf{u})=0$ is referred to as the nonlinearly preconditioned system. The latter is generated by applying Newton’s or inexact Newton’s method to a fixed point equation. The goal is for the nonlinearly preconditioned system to be solved faster than the original one; that is, to more quickly advance to the region of quadratic convergence for Newton’s method and reduce the number of exact or inexact Newton iterations.
@@
In the Additive Schwarz Preconditioned Inexact Newton (ASPIN) [7], the authors introduce a method, where the nonlinearly preconditioned system $\mathcal{F}(u)=0$ consists of more balanced/uniform nonlinearities. ASPIN involves the solution of local nonlinear problems, using similar restriction and extension operators that are used in the linear case. Then, a global linear problem is solved via inexact Newton. In theory, for difficult problems with intense nonlinearities, this technique can accelerate convergence when compared to a typical Newton or inexact Newton method. Additionally, it can aid in the occurrence of stagnation in Newton’s method and desensitize Newton to the initial guess. We remark as well the existence of the MSPIN method [26], where generally, the partitioning of degrees of freedom is based on field type (field-splitting) instead of the typical subdomain partitioning. A main difference between the methods is that MSPIN can be thought of as analogous to the Multiplicative Schwarz (MS) method for linear systems, a generally non-parallel method, while ASPIN is analogous to the Additive Schwarz method (ASM). With this, one can also think of ASPIN and MSPIN as varying from each other similarly to Jacobi and Gauss-Seidel methods. In MSPIN and ASPIN, inexact Newton solves are computed for the nonlinearly preconditioned system.
@@
The authors of [11] introduced SRASPEN, which is obtained similarly by applying Newton’s method to the fixed point equation resulting from the nonlinear SRAS (Substructured Restricted Additive Schwarz) method. Unlike the aforementioned methods, the SRASPEN method works on nonoverlapping subdomains; these nonoverlapping domain decomposition methods are often referred to as substructuring methods. Within the same spirit, the authors of [12] introduced DNPEN, a nonlinear preconditioning method which applies Newton’s method to the Dirichlet-Neumann (DN) method, which is a substructuring method based on matching fluxes along subdomain interfaces. Additionally, the authors of [27] propose an adaptive nonlinear preconditioning technique which “turns off” the nonlinear preconditioning for outer Newtons where it is not necessary.
@@
The aforementioned techniques can be considered as being left preconditioning methods, as they are solving a new tangential system $\mathcal{F}(\mathbf{u})=0$ by exact or inexact Newton. We also remark that there are a family of right nonlinear preconditioning techniques; these methods change the unknowns of the original system. We can give as an example the following works [22, 28, 23, 24, 25], where the authors introduced nonlinear preconditioners analogous to the Finite Element Tearing and Interconnecting (FETI), Finite Element Tearing and Interconnecting Dual-Primal (FETI-DP), and Balancing Domain Decomposition by Constraints (BDDC) methods. Additionally, the authors of [8] introduced the Nonlinear Elimination Preconditioned Inexact Newton (NEPIN), where “problem” components of the residual are implicitly eliminated and a resulting modified Newton direction is computed.
@@
As is the case for linear DD methods, two-level methods (methods which include a coarse correction/level), are often needed for scalability in nonlinear preconditioning methods. Particularly, the number of GMRES iterations for the global tangential solve will not be scalable without a coarse correction. The coarse component allows for global communication across all subdomains. In the original article introducing RASPEN [15], the authors proposed a Full Approximation Scheme (FAS) to apply the coarse level. For MSPIN, a multiplicative coarse correction counterpart was introduced. In [19], the author proposes a linear coarse level for ASPIN and in [17], the authors provide numerical evidence of scalability of two-level variants of ASPIN and RASPEN. In particular, they propose multiple approaches of adding the coarse space correction (both additively and multiplicatively) based on Galerkin projections, and a numerical comparison of the methods is reported for different types of coarse spaces.
@@
Our main contributions are as follows. In this paper, we rely on the work done in [4], in which we introduced an MsFEM coarse space for the linear Poisson problem posed on the perforated domains. Here, for nonlinear problems on perforated domains, we focus on left nonlinear preconditioning based on overlapping subdomains, coupled to techniques that use the coarse space proposed in [4] to provide scalability. We wish to stress that this article tackles the dual challenge of multiscale phenomena and double nonlinearity in the model problem; simultaneously addressing these points increases the complexity and significance of our approach. We propose the linear coarse space from [4] as a component of a two-level Restricted Additive Schwarz preconditioner for a linearized Newton system, which arises from the linearization of a nonlinear equation. Additionally, we use this coarse space as a component for multiple nonlinear methods, including a Two-level RASPEN method and a “Two-step” method originally introduced in [9] under the name NKS-RAS. While the coarse space of [4] was originally created based on the linear Poisson equation, we find that it is a fitting coarse space for nonlinear problems as well, and we use this in the later-mentioned two-level nonlinear preconditioning methods. The proposed methods also include the Anderson acceleration of a two-level method [31]. We provide a detailed complexity analysis of the proposed methods, discussing the benefits and downsides of each method. For these nonlinear methods applied to a time-dependent problem, we implement an adaptive local time-stepping method, which allows us to avoid a global time step reduction. In addition, this article serves as a numerical application of nonlinear preconditioning methods applied to nonlinear problems, containing multiple numerical examples with the main example being the doubly nonlinear Diffusive Wave model posed on complex domains containing multiple perforations.
@@
The remainder of the paper is structured as follows. Section 2 provides a detailed description of the finite-element/ finite-volume hybrid discretization used for the Diffusive Wave Equation. Section 3 describes the coarse space from [4], which will be used in the two-level methods described in the article. Section 4 summarizes domain decomposition methods for linear problems, presenting the well-known linear Additive Schwarz and Restricted Additive Schwarz methods. Section 5 describes the proposed methods in detail and discusses the cost/complexity of each proposed algorithm. Section 6 provides numerical results, with experiments based on a porous medium equation posed on both an L-shaped domain and a large urban domain, followed by the Diffusive Wave equation posed on a large urban domain. Section 7 concludes with a summary and description of future work.
2 Discretization of Diffusive Wave equation
@@
We now proceed with a discretization of the Diffusive Wave model (1). Let $T_{f}>0$ be the final flow simulation time and let $0=t_{0}<t_{1}<\ldots<t_{N_{T}}=T_{f}$ be a family of real numbers such that $\Delta t_{n}=t_{n+1}-t_{n}$.
@@
$\left\{\begin{array}[]{rll}\displaystyle\frac{u^{n+1}-u^{n}}{\Delta t_{n}}-\text{div}\left(c_{f}\kappa(u^{n+1},\nabla u^{n})\nabla u^{n+1}\right)&=&0\qquad\mbox{in}\qquad\Omega,\\ \displaystyle u^{n+1}&=&g\qquad\mbox{on}\qquad\partial\Omega\setminus\partial\Omega_{S},\\ \displaystyle c_{f}\kappa(u^{n+1},\nabla u^{n})\frac{\partial u^{n+1}}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad\partial\Omega\cap\partial\Omega_{S},\end{array}\right.$ (2)
@@
Let $(\mathbf{x}_{i})_{i=1}^{N_{\Omega}}$ denote the set of triangulation points; the set of point indices $\{1,2,...,N_{\Omega}\}$ is denoted by $\mathcal{N}$. We denote the nodal “hat” basis functions $(\eta^{\ell})_{\ell=1}^{N_{\Omega}}$. Furthermore, we define $u_{i}^{n+1}$ and $u_{\ell}^{n+1}$ as the nodal values of the discrete solution at nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{\ell}$ at $t_{n+1}$, respectively. Additionally, we set ${\cal N}_{T}=\{i\in{\cal N}\,\,|\,\,\mathbf{x}_{i}\in\overline{T}\}$ for any triangle $T\in{\cal T}$ ($T$ assumed to be open) and denote $\mathcal{T}_{i}$ as the subset of triangles connected to the node $i$ such that $\mathcal{T}_{i}=\{T\in\mathcal{T}\,\,|\,\,\mathbf{x}_{i}\in\overline{T}\}$. Furthermore, we denote by ${\cal N}_{i}$ the set of notes adjacent to node $i$ such that ${\cal N}_{i}=\bigcup_{T\in\mathcal{T}_{i}}{\cal N}_{T}.$ Finally, let ${\cal N}_{D}$ denote the set of Dirichlet nodes such that ${\cal N}_{D}=\{i\in{\cal N}\,\,|\,\,\mathbf{x}_{i}\in\partial\Omega\setminus\Omega_{S}\}$.
In order to enhance the stability of the numerical scheme, we modify the finite element scheme by means of discretization techniques from finite volume methods. In particular, we introduce mass lumping in the accumulation and upwinding in the diffusion term. This discretization technique is generally referred to as a Finite Volume- Finite Element (FV-FE) discretization [13], or it is often referred to as a Control Volume Finite Element method [10, 14].
@@
$\tau_{i\ell,T}^{n}=-c_{f}\Bigl{|}\nabla u_{h}^{n}|_{T}\Bigr{|}^{1-\gamma}\int_{T}\nabla\eta_{\ell}\cdot\nabla\eta_{i}\,{\rm d}\mathbf{x}\qquad\mbox{and}\qquad\tau_{i\ell}^{n}=\sum_{T\in{\cal T}_{i}\cap{\cal T}_{\ell}}\tau_{i\ell,T}^{n},$